OpenStudy (anonymous):

solve for x and what is the root? square root of 3x+1=7

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\sqrt{3x+1}=7\] square both sides and solve \[3x+1=49\] in two steps

OpenStudy (anonymous):

oh, if it is \[\sqrt{3x}+1=7\] then subtract 1 from both sides and get \[\sqrt{3x}=6\] then square both sides to get \[3x=36\] and solve in one step